Problem 5-20 CVP Applications: Break-Even Analysis; Cost Structure; Target Sales [LO5-1, LO5-3, LO5- 4, LO5-5, LO5-6, LO5-8] Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost. Last year, the company sold 46,000 of these balls, with the following results Sales (46,000 balls) Variable expenses Contribution margin Fixed expenses Net operating income $ 1,150,000 690,000 460,000 318,000 $ 142,000 Required: 1. Compute (a) last years CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last years sales level 2. Due to an increase in labor rates, the company estimates that next years variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next years CM ratio and the break-even point in balls? 3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $142,000, as last year? 4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs? 5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the companys new CM ratio and new break-even point in balls? 6. Refer to the data in (5) above a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $142,000, as last year? b. Assume the new plant is built and that next year the company manufactures and sells 46,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage Complete this question by entering your answers in the tabs below Req 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Compute (a) last years CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last years sales level. (Round Unit sales to break even to the nearest whole unit and other answers to 2 decimal places.) CM Ratio Unit sales to break even Degree of operating leverage ballsReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Due to an increase in labor rates, the company estimates that next years variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next years CM ratio and the break-even point in balls? (Round CM Ratio to 2 decimal places and Unit sales to break even to the nearest whole unit.) CM Ratio Unit sales to break even ballsReq 4 Req 5 Req 6A Req 6B Req 1 Req 2 Req 3 Refer to the data in Required (2). If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $142,000, as last year? (Round your answer to the nearest whole unit.) Number of ballsReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Refer again to the data in Required (2). The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs? (Round your answer to 2 decimal places.) Selling priceReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the companys new CM ratio and new break-even point in balls? (Round CM Ratio to 2 decimal places and Unit sales to break even to the nearest whole unit.) Show less CM Ratio Unit sales to break even ballsReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $142,000, as last year? (Round your answer to the nearest whole unit.) Number of ballsReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Assume the new plant is built and that next year the company manufactures and sells 46,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage. (Round Degree of operating leverage to 2 decimal places.) Northwood Company Contribution Income Statement Degree of operating leverage
Answer #1
Unit selling price 25 = 1150000/46000
Unit variable cost 15 = 690000/46000
1
CM ratio 40% =(25-15)/25
Units sales to break even 31800 =318000/(25-15)
Contribution margin 460000
Divide by Net operating income 142000
Degree of operating leverage 3.24
2
Revised variable expenses 18 = 15+3
CM ratio 28% =(25-18)/25
Units sales to break even 45429 =318000/(25-18)
3
Number of balls 65714 =(318000+142000)/(25-18)
4
Selling price 30 =18/(1-40%)
5
Revised variable expenses 9 =15*(1-40%)
Revised fixed cost 636000 =318000+318000
CM ratio 64% =(25-9)/25
Units sales to break even 39750 =636000/(25-9)
6a
Number of balls 48625 =(636000+142000)/(25-9)
b
Sales 1150000 =46000*25
Variable expenses 414000 =46000*9
Contribution margin 736000
Fixed expenses 636000
Net operating income 100000
Degree of operating leverage 7.36 =736000/100000